The volume of a certain metal varies directly with its mass. If a sample with mass 60.2 grams has a volume of 14.0 cubic centimeters, then what is the volume of a sample with mass 150.5 grams?
: The volume of a certain metal varies directly with its mass. If a sample with mass 60.2 grams has a volume of 14.0 cubic centimeters, then what is the volume of a sample with mass 150.5 grams?

Answer by ankor@dixie-net.com(4538)

The volume of a certain metal varies directly with its mass. If a sample with mass 60.2 grams has a volume of 14.0 cubic centimeters, then what is the volume of a sample with mass 150.5 grams?
:
Let x = vol of 150.5 gr
:
Use ratio:
x/14 = 150.5/60.2
Cross mult
60.2x = 14*150.5
:
x = 2107/60.2
x = 35 cu cm